A refrigerator is to maintain eatables kept inside at 9 c. If room temperature is 36 c, calculate the performance

Grade:Upto college level

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
Coefficient of performance of a refrigerator is given as
COP=\frac{T_{cold}}{T_{hot}-T_{cold}}
=282/(309-282)= 10.444
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty